In a coffee-cup calorimeter, 11.0-g sample of solid CaCℓ2 is dissolved in 125 g of water at 25.0 oC. The temperature in the calorimeter is measured to be 39.2 oC when the dissolution of CaCℓ2 is completed. Assuming that the specific heat of solution is equal to that of water, i.e., 4.184 J/g oC, calculate the heat of solution of CaCℓ2 in water, in kJ/mol.


final answer is :-29.6
Answer #1

Moles of CaCl2 = mass / molar mass = 11 / 111 = 0.0991 moles

Heat released = mass of water * temperature change * Specific heat = 125 * (39.2 - 25) * 4.184 = 7426.6 J = 7.427 KJ

Enthalpy of solution = Heat released / moles of caCl2 = -7.427 / 0.0991 = -74.94 KJ/mol